# SHuBERT: Self-Supervised Sign Language Representation Learning via Multi-Stream Cluster Prediction

### v1

@@ 4 Experiments and Results
Below we describe the experimental setup for self-supervised training of SHuBERT, followed by its adaptation as a foundation model for multiple sign language processing tasks: sign language translation (Sec. 4.2) isolated sign language recognition (Sec. 4.3), and phonological feature recognition (Sec. 4.4).
@@ 4.1 Pre-Training SHuBERT
For pre-training SHuBERT, we use the Youtube-ASL dataset [55] excluding the clips that intersect with the OpenASL dataset [50], so that we can use OpenASL to evaluate SHuBERT. To maintain the same training set size as the original YouTube-ASL dataset, we replaced the removed content with ASL videos from YouTube-SL-25 [54] that are not present in YouTube-ASL. Our final pre-training dataset comprises approximately 987 hours of ASL content.
To reduce computation, we downsample the videos by removing every other frame The average frame rate of the post-processed videos is 14.89 fps.
@@
We train a base model consisting of 12 transformer blocks, with each block having an embedding dimensionality of 768, feed-forward dimensionality of 3072, and 12 attention heads. The complete model contains 86M parameters. For each channel (face, left hand, right hand, and body posture), we use $k$-means on 10% of the data to create 256 discrete clusters that serve as prediction targets. Our masking strategy uses a span length of 3 frames (approximately 200ms), which is the average duration of fingerspelling a single letter in American Sign Language and is therefore roughly the smallest gesture length  [57]. We train the model for 400K steps using 8 NVIDIA A6000 (48GB) GPUs, with a total training time of approximately 7 days.
We optimize the model using Adam with a peak learning rate of $5\times 10^{-4}$, warming up for the first $8\%$ of updates followed by linear decay. Videos are batched to maintain efficiency while not exceeding 1,500 frames per GPU.
@@ 4.2 Sign Language Translation
For sign language translation, where the input sign language video is mapped to text in English, we use ByT5-Base [61] as a translation model to map from SHuBERT representations to written English, following prior work showing its strong performance on this task [52, 63]. We first extract video representations from SHuBERT and project them to ByT5’s input space through a simple linear layer. We train the combined model with a cross-entropy loss and label smoothing factor of 0.2. We use a two-phase training strategy. In the first phase, we train the translation system on the weakly labeled YouTube-ASL dataset (with OpenASL removed, as described above) for 250K steps. During this phase, we use the AdamW optimizer with a peak learning rate of $5\times 10^{-4}$ for ByT5 and a reduced learning rate of $5\times 10^{-5}$ for SHuBERT parameters (when fine-tuned). The learning rate follows a cosine schedule with 10K warmup steps. We use a batch size of 2 utterances per GPU with gradient accumulation over 8 steps and use weight decay of 0.1 for regularization.
In the second phase, we fine-tune on two target benchmark datasets (How2Sign [13] and OpenASL [50]) for 50K steps, using a lower learning rate of $10^{-4}$ and 5K warmup steps. We also evaluate in a zero-shot setting, without any additional fine-tuning, on a third dataset (for which no training data exists), FLEURS-ASL [53]. As often done when using pre-trained speech encoders [58], we use a learned weighted sum of features from all SHuBERT layers rather than using a single layer’s output. During decoding, we use beam search with a beam width of 5, and a maximum sequence length of 384 tokens.
Table 1: Translation results on How2Sign. SSL: self-supervised learning (yes/no); PT: pre-training data.
@@
Tanzer [52] $\times$ $\sim$2,800 18.1 50.8
Zhang et al. [63] $\times$ $\sim$6,600 21.1 55.7
@@
SSVP [44] $\checkmark$ 1,054 15.5 49.6
Uthus et al. [55] $\times$ 984 12.4 46.6
Tanzer and Zhang [54] $\times$ 3,207 15.4 47.9
SignMusketeers [19] $\checkmark$ 984 14.3 -
Ours $\checkmark$ 984 16.2 49.9
Table 2: Translation results on OpenASL
@@
Publicly available data
OpenASL[50] $\times$ - 6.7 31.1
GloFE-VN [33] $\times$ - 7.1 36.7
C2RL [10] $\times$ - 13.2 -
Ours $\checkmark$ 984 23.2 60.6
Table 3: Translation results on FLEURS-ASL
[…]
Private data
Tanzer [52] $\times$ $\sim$2,800 5.8 45.4
Publicly available data
Tanzer [53] $\times$ 984 3.9 38.3
Tanzer and Zhang [54] $\times$ 3,207 4.4 40.1
Ours $\checkmark$ 984 4.7 41.2
We evaluate the final model on the three benchmark test sets. Tab. 1, Tab. 2, and Tab. 3 show translation results on How2Sign, OpenASL, and FLEURS-ASL, respectively. We evaluate using the standard BLEU[39, 40] and BLEURT[47] translation metrics.
In all cases, our results using SHuBERT improve over the best prior published results using publicly available training data. On How2Sign, SHuBERT achieves a +0.7 BLEU/+0.3 BLEURT improvement over the prior state-of-the-art result (using public data), belonging to SSVP which used slighty more pre-training data (1,054 vs. 984 hours, as they included the training data of How2Sign in pre-training). Better published results exist (as shown in Tab. 1), but they use private fine-tuning datasets so we cannot reproduce their settings nor compare to them meaningfully.
In the case of OpenASL, SHuBERT’s improvement over the best prior result is particularly large (+10.0 BLEU/+23.9 BLEURT). This may be attributable to pre-training on similar-domain (native, natural rather than interpreted signing) data in YouTube-ASL (but, as previously mentioned, not any of the same data). The distinction between interpreted signing in a constrained visual environment (as in How2Sign and FLEURS-ASL) and natural signing in a less constrained environment (as in OpenASL and much of YouTube-ASL) is an important one that has not received sufficient attention in the sign language understanding literature, and is worth exploring further in future work. The properties of the visual environment affect the difficulty of the task, and natural signing has different characteristics from those of interpreted sign language, often described as “translationese” [16, 29].
We also note that most previous approaches use additional enhancement techniques shown to contribute to their final scores, such as an additional contrastive learning with labelled data [44], joint training with text machine translation [63], and multi-tasking with random dynamic clips [53]. In contrast, our setup uses none of those. It is possible that incorporating such techniques into our framework will further improve scores, but we leave this for future work.
4.3 Isolated Sign Language Recognition
For isolated sign language recognition (ISLR), the task of classifying a short video of a single sign, we compare a range of approaches for using SHuBERT, including learning LoRA adapters [21], fully fine-tuning, and joint multi-task training on multiple ISLR datasets. Unless otherwise specified, for all experiments, we train for 125 epochs with a batch size of 128 and perform early stopping according to validation results. We use an Adam optimizer [28] with a learning rate of $10^{-4}$ and a weight decay of $10^{-4}$. For adapting SHuBERT to classification tasks, we first average SHuBERT representations across the time dimension and add a batch-norm layer followed by a linear layer as the classification head.
For LoRA training, we learn a rank-1 LoRA module for each linear layer in SHuBERT while keeping all the other weights frozen, increasing by only 0.197% the number of parameters. In addition to the aforementioned hyperparameters, we reduce the learning rate of the LoRA modules to 1/10 of the classification head’s and also use a 0.1 label smoothing for further regularization.
For the case of full fine-tuning of SHuBERT, the training protocol is largely the same as for LoRA training, except that we re-initialize the last layer of SHuBERT which we find to give a small performance improvement.
Finally, inspired by prior work such as [55] and [44], where training on larger datasets greatly enhances the performance on smaller ones, we fine-tune SHuBERT jointly on three ISLR datasets: ASL Citizen [11], WLASL2000 [32, 37] (cleaned version) and Sem-Lex [26]. Since these datasets have different labels, we frame the learning as multi-task training; each task has its own classifier, with the fine-tuned SHuBERT encoder shared across tasks. The losses from the three datasets are weighted equally. The hyperparameters are the same as in full fine-tuning except for the batch size, which is reduced to 100 due to GPU memory constraints.
Our ISLR results on ASL Citizen, WLASL2000 (clean) and Sem-Lex are shown in Tab. 4, Tab. 5 and Tab. 6, respectively. We report Recall at 1, 5 and 10. We achieve state-of-the-art performance on all datasets except WLASL2000 (clean). [^2] Note: As described in the caption of Tab. 6, the Sem-Lex results in [26] are not comparable with other methods, including ours. As shown in all three tables, joint training outperforms the other two training strategies on all datasets we experiment on, with a margin of 3.8% on average. Rank-1 LoRA, although it is the worst performer of our results on two of the three datasets, still outperforms previous SOTA results on ASL Citizen and Sem-Lex. For WLASL2000 (clean), Ours (joint) yields competitive results that are nearly on par with MaskFeat, a model trained specifically for ISLR.
Table 4: ISLR accuracy results on the ASL Citizen dataset.
Model Rec@1$\uparrow$ Rec@5$\uparrow$ Rec@10$\uparrow$
ST-GCN[11] 0.60 0.82 0.88
SignCLIP[25] 0.60 0.84 0.89
I3D[11] 0.63 0.86 0.91
Ours (rank=1 LoRA) 0.65 0.87 0.91
Ours (fine-tune) 0.65 0.87 0.91
Ours (joint) 0.68 0.90 0.93
Table 5: ISLR accuracy results on the WLASL2000 (clean) dataset.
Model Rec@1$\uparrow$ Rec@5$\uparrow$ Rec@10$\uparrow$
SVT[45] 0.19 - -
BEVT[45] 0.61 - -
VideoMAE[45] 0.69 - -
MaskFeat[45] 0.79 - -
Ours (rank=1 LoRA) 0.68 0.89 0.94
Ours (fine-tune) 0.71 0.92 0.95
Ours (joint) 0.76 0.94 0.96
@@ 4.3 Isolated Sign Language Recognition
Table 6: ISLR accuracy results on the Sem-Lex dataset. The result marked with an asterisk () is not directly comparable to Ours as it is for a reduced (and easier) test set, as mentioned in [25]. Additionally, the dataset version released by [26] has a significant fraction of videos missing.
Model Rec@1$\uparrow$ Rec@5$\uparrow$ Rec@10$\uparrow$
Sem-Lex[26] $0.69^{*}$
SignCLIP[25] 0.30 0.48 0.55
Ours (rank=1 LoRA) 0.54 0.74 0.80
Ours (fine-tune) 0.53 0.73 0.79
Ours (joint) 0.58 0.78 0.83
4.4 Phonological Feature Recognition
Table 7: Phonological feature recognition accuracy results of two datasets, Sem-Lex and ASL Citizen.

### v3

@@ 4 Experiments and Results
In this section, we describe the experimental setup for self-supervised training of SHuBERT, followed by its adaptation as a foundation model for multiple sign language processing tasks: sign language translation (Sec. 4.2), isolated sign language recognition (Sec. 4.3) and fingerspelling detection (Sec. 4.4). We also apply SHuBERT to phonological feature recognition, as a baseline for future work (see Appendix B).
@@ 4.1 Pre-Training SHuBERT
For pre-training SHuBERT, we use the YouTube-ASL dataset Uthus et al. (2023). Note that we excluded the clips that intersect with the OpenASL dataset Shi et al. (2022) to evaluate SHuBERT on OpenASL. To maintain the same training set size as the original YouTube-ASL dataset, we replaced the removed content with ASL videos from YouTube-SL-25 Tanzer and Zhang (2024) that are not present in YouTube-ASL. Our final pre-training dataset comprises approximately 984 hours of ASL content. [^4] Note that YouTube-ASL encompasses parts of several datasets used in other work, including 72.4% of the OpenASL test set and 38.2% of the MSASL Joze and Koller (2018) test set. We do not compare to other work on MSASL for this reason.
To reduce computation, we downsample the videos by removing every other frame Uthus et al. (2023). The average frame rate of the post-processed videos is 14.89 fps.
@@
We train a base model consisting of 12 transformer blocks, with each block having an embedding dimensionality of 768, feed-forward dimensionality of 3072, and 12 attention heads. The complete model contains 86M parameters. For each channel (face, left hand, right hand, and body pose), we use $k$-means on 10% of the data to create 256 discrete clusters that serve as prediction targets. Figs. 5, 6, 7 and 8 in the Appendix provide examples of images in several clusters. Our masking strategy uses a span length of 3 frames (approximately 200ms), which is roughly the average duration of fingerspelling a single letter in ASL and is therefore roughly the smallest gesture length  Hanson (1982). We train the model for 400K steps using 8 NVIDIA A6000 (48GB) GPUs, with a total training time of approximately 7 days.
We optimize the model using Adam with a peak learning rate of $5\times 10^{-4}$, warming up for the first $8\%$ of updates followed by linear decay. We batch videos to maintain efficiency while not exceeding 1,500 frames per GPU.
@@ 4.2 Sign Language Translation
For sign language translation, where the input sign language video is mapped to text in English, we use ByT5-Base Xue et al. (2022), pre-trained on a large corpus of unlabeled multilingual text data, as a translation model to map from SHuBERT representations to written English, following prior work showing its strong performance on this task Tanzer and Zhang (2024); Zhang et al. (2024). We first extract video representations from SHuBERT and project them to ByT5’s input space through a linear layer. We train the combined model (SHuBERT+projection layer+ByT5) with a cross-entropy loss and label smoothing factor of 0.2.
Similarly to Uthus et al. (2023) and Rust et al. (2024), we use a two-phase training strategy. In the first phase, we train the translation system on the weakly labeled YouTube-ASL dataset (with OpenASL removed, as described above) for 250K steps. During this phase, we use the AdamW optimizer with a peak learning rate of $5\times 10^{-4}$ for ByT5 and a reduced learning rate of $5\times 10^{-5}$ for SHuBERT parameters (when fine-tuned). The learning rate follows a cosine schedule with 10K warmup steps. We use a batch size of 2 utterances per GPU with gradient accumulation over 8 steps and use weight decay of 0.1.
In the second phase, we fine-tune on two target benchmark datasets (How2Sign Duarte et al. (2021) and OpenASL Shi et al. (2022)) for 50K steps, using a lower learning rate of $10^{-4}$ and 5K warmup steps. We also evaluate in a zero-shot setting, without any additional fine-tuning, on a third dataset (for which no training data exists), FLEURS-ASL Tanzer (2024b). We use a learned weighted sum of features from all SHuBERT layers rather than using a single layer’s output, as is commonly done when using speech representations such as HuBERT Yang et al. (2021). During decoding, we use beam search with a beam width of 5 and a maximum sequence length of 384 tokens.
Method SSL PT data (hrs) How2Sign OpenASL FLEURS-ASL
@@
Tanzer (2024a) $\times$ $\sim$2,800 18.1 50.8 - - 5.8 45.4
Zhang et al. (2024) $\times$ $\sim$6,600 21.1 55.7 - - - -
@@
SSVP Rust et al. (2024) $\checkmark$ 1,054 15.5 49.6 - - - -
Uthus et al. (2023) $\times$ 984 12.4 46.6 - - - -
Tanzer and Zhang (2024) $\times$ 3,207 15.4 47.9 - - 4.4 40.1
SM Gueuwou et al. (2025) $\checkmark$ 984 14.3 - - - - -
VAP Jiao et al. (2024) $\times$ - 12.9 - 21.2 - - -
Uni-Sign Li et al. (2025) $\times$ 984 14.9 49.4 23.1* 60.4* - -
OpenASL Shi et al. (2022) $\times$ - - - 6.7 31.1 - -
GloFE-VN Lin et al. (2023) $\times$ - - - 7.1 36.7 - -
C2RL Chen et al. (2025) $\times$ - - - 13.2 - - -
Tanzer (2024b) $\times$ 984 - - - - 3.9 38.3
Ours $\checkmark$ 984 16.2 49.9 23.2 60.6 4.7 41.2
Table 1: Translation results on the How2Sign, OpenASL, and FLEURS-ASL test sets. SSL: self-supervised learning (yes/no); PT: pre-training. * Uni-Sign is pre-trained on YouTube-ASL, which contains >72% of the OpenASL test samples, so we do not consider the results on OpenASL to be directly comparable to ours.
We evaluate the final model on the three benchmark test sets of How2Sign, OpenASL, and FLEURS-ASL, using the standard BLEU Papineni et al. (2002); Post (2018) [^5] SacreBLEUversionsignature:BLEU+c.mixed+#.1+s.exp+tok.13a+v.1.4.1. and BLEURT Sellam et al. (2020) translation metrics, as shown in Tab. 1 (see also example translations in Tabs. 11, 12 and 13).
In all cases, our results using SHuBERT improve over the best prior published results using publicly available training data. On How2Sign, SHuBERT improves by +0.7 BLEU/+0.3 BLEURT over the prior state-of-the-art result (using public data) of SSV-SLT Rust et al. (2024), which used slighty more pre-training data (1,054 vs. 984 hours) that included the training data of How2Sign. Better published results exist (as shown in Tab. 1), but they rely on private fine-tuning datasets so we cannot reproduce their settings nor compare to them meaningfully.
In the case of OpenASL, SHuBERT’s improvement over the best prior result is +2.0 BLEU. [^6] Uni-Sign Li et al. (2025) reports a similar BLEU score of 23.1 on OpenASL, but Uni-Sign is pre-trained on YouTube-ASL. Therefore, most of the test set is included in Uni-Sign’s pre-training data, so we do not consider the results directly comparable. This larger improvement may be attributable to pre-training on similar-domain (generally native, natural rather than interpreted signing) data in YouTube-ASL (but, as previously mentioned, none of the same data). The distinction between interpreted signing in a constrained visual environment (as in How2Sign and FLEURS-ASL) and natural signing in a less constrained environment (as in OpenASL and much of YouTube-ASL) is an important one that has not received sufficient attention, and is worth exploring further in future work. The properties of the visual environment affect the difficulty of the task, and natural signing has different characteristics from those of interpreted sign language Desai et al. (2024c).
We also note that some previous approaches use additional techniques such as auxiliary losses that contribute to their final results, such as additional contrastive learning with labelled data Rust et al. (2024), joint training with text machine translation Zhang et al. (2024), and multi-tasking with random dynamic clips from an original video Tanzer (2024b). It is possible that incorporating such techniques into our framework will further improve performance, but we leave this for future work.
Ablations We conduct several ablation studies to validate our design choices and analyze SHuBERT’s behavior. These studies examine: (1) the impact of different masking strategies during pre-training, where random masking proves most effective based on BLEURT scores; (2) the importance of pre-training data scale, showing the clear benefit of using the full pre-training dataset; (3) the contribution of different architectural components, demonstrating that a weighted combination of layers significantly improves translation performance; and (4) the effects of fine-tuning versus keeping SHuBERT frozen during translation training, with fine-tuning providing moderate gains. The strong performance of the frozen, layer-weighted SHuBERT suggests that it is a promising approach for low-resource settings where parallel data may be limited. Detailed results and analysis of these ablations can be found in Appendix A.
4.3 Isolated Sign Language Recognition
Method #Params ASL Citizen Sem-Lex WLASL2000
@@ 4.3 Isolated Sign Language Recognition
ST-GCN Desai et al. (2024a) 0.45M 0.60 0.82 0.88 - - - - -
SignCLIP Jiang et al. (2024) 217M 0.60 0.84 0.89 0.30 0.48 0.55 - -
I3D Desai et al. (2024a) 25M 0.63 0.86 0.91 - - - - -
Sem-Lex Kezar et al. (2023) 0.45M - - - $0.69^{*}$
SignBERT Hu et al. (2021) - - - - - - - 39.40 36.74
SignBERT+ Hu et al. (2023) - - - - - - - 48.85 46.37
MSLU Zhou et al. (2024) - - - - - - - 56.29 53.29
NLA-SLR Zuo et al. (2023) - - - - - - - 61.05 58.05
Uni-Sign Li et al. (2025) 580M - - - - - - 63.52 61.32
Ours (rank=1 LoRA) 0.17M 0.65 0.87 0.91 0.54 0.74 0.80 60.90 58.01
@@
Table 2: ISLR results on the ASL Citizen, Sem-Lex, and WLASL2000 test sets. Note: For Sem-Lex, the result marked with an asterisk () is not directly comparable to Ours as it is for a reduced (and easier) test set, as mentioned in Jiang et al. (2024). Additionally, the dataset version released by Kezar et al. (2023) has a significant fraction of videos missing. For WLASL2000, evaluation metrics are per-instance (P-I) and per-class (P-C) Top-1 accuracy.
For isolated sign language recognition (ISLR), the task of classifying a short video of a single sign, we include results for SHuBERT adapted with LoRA adapters Hu et al. (2022). Unless otherwise specified, for all experiments, we train for 125 epochs with a batch size of 128 and perform early stopping according to validation results (R@1/P-I). We use an Adam optimizer Kingma (2014) with a learning rate of $10^{-4}$ and weight decay of $10^{-4}$. For classification tasks, we first average SHuBERT representations across the time dimension and add a batch-norm layer followed by a linear layer as the classification head.
For LoRA training, we learn a rank-1 LoRA module for each linear layer in SHuBERT while keeping all the other weights frozen, resulting in training only 0.2% of the number of parameters of the original model. In addition to the aforementioned hyperparameters, we reduce the learning rate of the LoRA modules to 1/10 of the classification head’s and also use 0.1 label smoothing.
Our ISLR results on ASL Citizen, WLASL2000 (original), and Sem-Lex are shown in Tab. 2. We note that we do not report on MSASL, as done in some of the prior work, because of the aforementioned overlap between its test set and YouTube-ASL. Following prior work, we report Recall at 1, 5 and 10, unless stated otherwise. [^7] Note that some prior work reports ISLR results in terms of accuracy, which is equivalent to Recall at 1. We achieve state-of-the-art performance on all datasets except WLASL2000 where Uni-Sign has a better result. We note that Uni-Sign fine-tunes 3,000 times more paramters than ours. [^8] Note: As described in the caption of Tab. 2, the Sem-Lex results in Kezar et al. (2023) are not comparable with other methods, including ours.
4.4 Fingerspelling detection
Method SSL Mean IoU $\uparrow$
Contrastive Learning Yin et al. (2024) $\times$ 0.28
SHuBERT (Ours) $\checkmark$ 0.40
Table 3: Fingerspelling detection on ASL-Stem Wiki Yin et al. (2024).
Finally, we evaluate SHuBERT on the task of fingerspelling detection on the ASL-Stem-Wiki dataset Yin et al. (2024). Given a sign language video input $v$, which consists of an ordered sequence of frames $\{v_{1},v_{2},\ldots,v_{n}\}$, the task is to identify all segments containing fingerspelling. The output is represented as a set $F$ of frame intervals: $F=\{[s_{1},e_{1}],[s_{2},e_{2}],\ldots,[s_{k},e_{k}]\}$, where each interval $[s_{i},e_{i}]$ represents the start and end frames of a fingerspelling sequence, such that frames $v_{s_{i}}$ through $v_{e_{i}}$ contain fingerspelling. We follow the original ASL-Stem-Wiki evaluation pipeline (cross-validation) and evaluation metric (intersection over union, or IoU). The results are shown in Tab. 3. We see that by simply fine-tuning SHuBERT for this task, we increase the IoU by 42% compared to the previous state of the art method, which pre-trains and fine-tunes on the same dataset.
5 Conclusion
SHuBERT, our proposed self-supervised approach for learning sign language video representations, yields a transformer encoder that maps from multiple feature streams (face and hand appearance and upper body pose) to a stream of per-frame contextual representations. A single base SHuBERT model, when adapted to a range of sign language processing benchmarks including both translation and isolated sign recognition, achieves strong performance on all of them and almost always improves over the prior state of the art. SHuBERT is trained on public data and is publicly available. [^9] http://shubert.pals.ttic.edu Based on its strong performance on the tasks studied here, we expect that SHuBERT can serve as a base model for a broad range of sign language processing tasks.
Limitations
Our work has several limitations. First, although the results are competitive with or outperform prior work, the absolute performance is still quite poor. Neither our model nor others can replace human interpreters for broad-domain sign language translation. Second, our training data volume is significantly smaller than that of typical self-supervised speech and text models. We cannot say with certainty how our model would scale up to much larger datasets. Third, we have not carefully studied potential sources of bias in the model. From our qualitative visual inspection of images and their corresponding clusters (Figs. 5, 6, 7 and 8 in the Appendix), we observe that semantic properties appear to take precedence over attributes like skin color, gender, or eyewear. While these preliminary observations are encouraging, a more thorough investigation of potential biases would be valuable future work. Finally, the current scope of our work is limited to American Sign Language. Although sign languages use the same channels and share many elements, we do not know how well our model would generalize to other languages. Future work could address these limitations by expanding the training dataset and training on data from other sign languages.
Acknowledgment
Appendix A Ablations on SHuBERT
Masking Strategies
Masking strategy BLEU-1 BLEU BLEURT
Channel masking 14.5 2.6 29.9
Time masking 15.1 2.3 31.2
Random masking 15.4 2.2 31.4
Tab. 4 compares translation performance when using each of the three masking strategies (Fig. 4). We observe different signals from different metrics. While the BLEU scores suggest channel masking to be the best, random masking produces better BLEURT scores. Contradicting signals between different evaluation metrics for sign language translation has also been observed in prior work Zhang et al. (2024). We chose to priotize BLEURT, as it generally has better alignment with human judgements of translation quality Freitag et al. (2022), and therefore use random masking in all of our other experiments.
Data Scaling Behavior of SHuBERT
Hours of pre-train data BLEU-1 BLEU BLEURT
984 15.4 2.2 31.4
98 12.7 0.7 29.1
Table 5: Impact of pre-training data size on SHuBERT’s performance.
In addition to the full pre-training dataset, we also train SHuBERT from scratch on a randomly selected 10% of the full data. In Tab. 5 we see that there is a noticeable drop in performance in BLEU and BLEURT, suggesting that data size is important. With multiple larger datasets now available—BOBSL Albanie et al. (2021), JWSign Gueuwou et al. (2023) and YouTube-SL-25 Tanzer and Zhang (2024) contain approximately 1500, 2500, and 3200 hours of data respectively—we expect that expanding the SHuBERT pre-training data may further improve performance. In addition, these larger datasets also include more language diversity, which may also improve performance and/or applicability to additional languages.
Isolating SHuBERT’s Impact on Performance
Layer of SHuBERT BLEU-1 BLEU BLEURT
None 15.3 2.5 31.6
Last Layer 21.4 4.7 35.0
Weighted Sum 29.3 7.1 39.5
To understand SHuBERT’s impact on translation performance, we conduct three experiments, shown in Tab. 6. In our baseline experiment (“None"), we directly feed the projected 4-channel features (face, left hand, right hand, body pose) to the ByT5 translation model, bypassing SHuBERT entirely, resulting in fairly poor performance. When we instead pass these features through a frozen pre-trained SHuBERT (400k steps, random masking) and use its final layer’s output (“Last Layer"), we see significant improvement. Finally, computing a learned weighted combination of all SHuBERT layers (“Weighted Sum") further improves performance. These results demonstrate that each component of SHuBERT contribute to translation performance.
Frozen vs. Fine-Tuned SHuBERT
Fine-tune SHuBERT BLEU-1 BLEU BLEURT
$\times$ 21.4 4.7 35.0
$\checkmark$ 30.0 7.5 39.9
We also investigate the impact of fine-tuning during translation training, as shown in Tab. 7. We compare two scenarios: using a frozen SHuBERT (400k steps, random masking) and only fine-tuning ByT5, versus fine-tuning both SHuBERT (from the same base model) and ByT5 together. Both scenarios use features from SHuBERT’s last layer. Fine-tuning SHuBERT leads to substantial improvements compared to keeping it frozen, when using the final layer. However, referring back to Tab. 6, the relatively small difference between the fine-tuned performance and that of the frozen and layer-weighted SHuBERT is noteworthy. This observation is promising for low-resource sign languages, where we may have plentiful unlabeled video data for pre-training, but very limited parallel data for translation training.
Stream contributions in ASL-to-English translation
Streams Used BLEU $\uparrow$
Face only 0.6
Hands only 0.2
Upper body only 0.8
Hands + Upper Body 2.1
All streams 2.4
Table 8: Stream contributions in ASL-to-English translation.
To quantify the contribution of each input stream, we conduct a translation experiment with the How2Sign dataset (without pre-training), feeding the concatenation of the multiple streams directly to a language model (T5). The resulting BLEU scores are shown in Tab. 8, showing that all of the streams contribute to translation. These results should be interpreted as assessing the true relative importance of each stream, however, since this is a very small-scale experiment.
Appendix B Phonological Feature Recognition